OpenStudy (anonymous):

Which of the following are true for E = mc2? Answer implies that a small change in mass produces a great deal of energy mainly used for chemical reactions implies that a small change is mass produces a small amount of energy

OpenStudy (abhisar):

As in your previous questions we saw that 1 kg mass can produce an energy equal to \(\sf 9\times 10^{16}J\) which is a farely large amount of energy. So option A is correct.

OpenStudy (anonymous):

thats what i was thinking, look i really appreciate all the help you have done and putting up with me

OpenStudy (abhisar):

Happy to help ! \(\Huge{\overset{\frown}{\normalsize \left( \begin{matrix} \Large\cdot \quad \cdot\\ \cdot\\ \huge \smile \end{matrix} \right)}}\normalsize \\ \;/\quad \;\;\quad \backslash\)